OpenStudy (anonymous):

integrate by partial fractions, not sure,. i want to integrate 1/((x^2*(x^2-15)) should i do partial fraction 1/((x^2*(x^2-15)) = A / x + B/x^2 + C/(x-sqrt 15) + D/(x+sqrt 15) , or 1/((x^2*(x^2-15)) = A / x + B/x^2 + Cx+d / ( x^2 - 15)

OpenStudy (anonymous):

why don't you type in math. will help in viewing easily.

OpenStudy (anonymous):

the second one is the right way. How do you hope getting radicals will help the partial fraction??(Indeed, we use partial fractions for simplification!!!)

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

whats wrong with partial fractions with radicals?

OpenStudy (anonymous):

i dont see any problem with the first way

OpenStudy (anonymous):

no problems but we avoid it as calculations get messy.

OpenStudy (anonymous):

my computer does the partial fractions integral 1/(x^2 (x^2-15)) = -1/(15x^2) + 1/(15(x^2-15))

OpenStudy (anonymous):

yah...that "seems" easy :) you can always take x^2=y during partial fracs.. then u get 1/(y(y-15)) which can be easily split into fracs.

OpenStudy (anonymous):

that doesnt disobey partial fractions rule

OpenStudy (anonymous):

i thought it was , it had to be linear factors and irreducibles

OpenStudy (anonymous):

like A/x + B/x^2 + C / (x-sqrt 15) + D/(x+sqrt 15)

OpenStudy (anonymous):

See...there are no so-called rules...at least not for me :) if we take radicals the math is correct but its unconventional

OpenStudy (anonymous):

cx+d is supposed to be only for irreducibles, i thought

OpenStudy (anonymous):

like integral 1/x(x^2+1) = A/x + bx+c / x^2 + 1

OpenStudy (anonymous):

so youre saying 1/x(x^2-1) = A/x + (bx+c) /( x^2 - 1)

OpenStudy (anonymous):

not only for irreducibles,

OpenStudy (anonymous):

im pretty sure thats what the theorem states

OpenStudy (anonymous):

maybe the theorem u saw said that we cx+d for irreducibles but it did not say that we cant use cx+d for non-irreducibles!!

OpenStudy (anonymous):

we use cx+d for irreducibles but we can also use it like (bx+c)/(x^2-1). Here the math is correct. But splitting like A/(x+1)+B/(x-1) is more advantageous, as it gives a good form whose integral is known.

OpenStudy (anonymous):

for example, this partial fractions fails 1/(x^2 (x-5) ) = A/x^2 + B/(x-5) , this fails

OpenStudy (anonymous):

oh because of the A/x + B/x^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

ok so 1/[(x-a)^n (x^2 +bx +c)^n] = k1/(x-a) + k2/(x-a)^2 + ... kn/(x-a)^n + m1x+p1/(x^2 +bx +c) +...

OpenStudy (anonymous):

your upto kn is ok. but from m1,p1,...its getting complicated.

OpenStudy (anonymous):

yeah im running out of letters

OpenStudy (anonymous):

no. there will be cx+d term above and the denominator's power will increase from 1 to n

OpenStudy (anonymous):

1/[(x-a)^n (x^2 +bx +c)^n] = A1/(x-a) + A2/(x-a)^2 + ... An/(x-a)^n + B1x+C1/(x^2 +bx +c) + (B2x +C2)/(x^2 +bx + c)^2... +

OpenStudy (anonymous):

1/[(x-a)^n (x^2 +bx +c)^n] = A1/(x-a) + A2/(x-a)^2 + ... An/(x-a)^n + B1x+C1/(x^2 +bx +c) + (B2x +C2)/(x^2 +bx + c)^2... +(Bnx +Cn) / (x^2 +bx +c)^n

OpenStudy (anonymous):

your c1 and b1 terms are wrong

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

there won't be any b1x ... also instead of c1 write it as b1x+c1

OpenStudy (anonymous):

1/[(x-a)^n (x^2 +bx +c)^n] = A1/(x-a) + A2/(x-a)^2 + ... An/(x-a)^n + (B1x+C1)/(x^2 +bx +c) + (B2x +C2)/(x^2 +bx + c)^2... +(Bnx +Cn) / (x^2 +bx +c)^n

OpenStudy (anonymous):

yups. thats correct....:)

OpenStudy (anonymous):

so thats fine?

OpenStudy (anonymous):

oh i left out the parenthesees

OpenStudy (anonymous):

ok so thats fine for partial fractions , and x^2 +bx + c doesnt necessarily have to be irreducible,

OpenStudy (anonymous):

see..make sure that you know that WE USE cx+d in numerator FOR IRREDUCIBLES (in other words polynomials have only complex roots). WE COULD use cx+d also above REDUCIBLES since it is mathematically correct but don't use it as one of the partial fractions is not in its lowest degrees (i.e. cx+d over a reducible is NOT recommended) But if such a use simplifies a problem at hand then why not use it? Understand me?

OpenStudy (anonymous):

a reducible quadratic you mean

OpenStudy (anonymous):

reducible into linear factors

OpenStudy (anonymous):

yups yups a QUADRATIC can only have cx+d above it. yes reducible into linear REAL factors. Nobody wants to have complex factors here!

OpenStudy (anonymous):

ok, so then we could do 1/(x^2 (x^2 - 15)) , we could reduce it

OpenStudy (anonymous):

im doing 1 / ( x ( x^2 + 2x + 1) ) the latter is reducible, but im not going to . = AX+b / ( x^2 + 2x + 1) + C / x

OpenStudy (anonymous):

Mathematically permissible. But not recommended.

OpenStudy (anonymous):

ok lets try this one 1/ ( x (x^2+2x+1)) integrate that using parfrac

OpenStudy (anonymous):

int 1/ ( x (x^2+2x+1))= (AX+b) / ( x^2 + 2x + 1) + C / x

OpenStudy (anonymous):

in part fracs.. its (1/x)+((-x-2)/(x^2+2x+1)) the first term gives log|x| and the second make substitution x^2+2x=y

OpenStudy (anonymous):

right, or we could have factored x^2 + 2x + 1

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

wait thats wrong

OpenStudy (anonymous):

whats that substitution ?

OpenStudy (anonymous):

yupss....the subs wrong...

OpenStudy (anonymous):

we can do it however

OpenStudy (anonymous):

oh by the way, partial fraction doesnt help us with say 1/ (x^(3/2) (x-15))

OpenStudy (anonymous):

no.. since the denominator is not a POLYNOMIAL

OpenStudy (anonymous):

, ahh

OpenStudy (anonymous):

both numer. and denom. must be polynomials

OpenStudy (anonymous):

wait is that a precondition?

OpenStudy (anonymous):

you could have 1/(x*(x^2-sqrt 3)) for example

OpenStudy (anonymous):

YES partial fractions can help ONLY TO RATIONAL FUNCTIONS i.e. both numeratoer and denominator polynomials with denominator not being the zero polynomial

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just clearing that up

OpenStudy (anonymous):

by the way what's your age?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

why do i sound dumb

OpenStudy (anonymous):

hahha..joking? i hope u understood.. ask for doubts...

OpenStudy (anonymous):

i am 17

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

jeez, i thought you were older. you are wise beyond your years :)

OpenStudy (anonymous):

are you a mathematician? or math-related worker? u named urself cantorset...

OpenStudy (anonymous):

yes im a math tutor

OpenStudy (anonymous):

for a living, i was stumped because the student didnt want to reduce a quadratic

OpenStudy (anonymous):

and i thought, hey the theorem says so and so

OpenStudy (anonymous):

it was reducible into irrational factors

OpenStudy (anonymous):

the question u asked me? - yes it can be solved both ways

OpenStudy (anonymous):

see this problem we just did its a lot easier if we reduce the quadratic

OpenStudy (anonymous):

integral (1/(x(x^2+2x+1))

OpenStudy (anonymous):

yes its lot easier if we reduce quad... but i argued just because i wanted to show that cx+d hold good for reducible quadratics also..

OpenStudy (anonymous):

right so i guess it depends on the problem

OpenStudy (anonymous):

well its good to know the power of this theorem

OpenStudy (anonymous):

which area of math are you specialized in?

OpenStudy (anonymous):

i tutor mostly calculus ,

OpenStudy (anonymous):

wow..thats awesome..i love calculus..but know little..struggling still with single variable calculus

OpenStudy (anonymous):

youre amazing :)

OpenStudy (anonymous):

well if you need help, give me a buzz

OpenStudy (anonymous):

ok..thank you.

OpenStudy (anonymous):

so yeah, this one is tougher to integrate , without reducing the quadratic

OpenStudy (anonymous):

integral (1/x)+((-x-2)/(x^2+2x+1))

OpenStudy (anonymous):

yups.

OpenStudy (anonymous):

so the original question integral 1/(x^2 (x^2-15))

OpenStudy (anonymous):

then reducing the quadratic is worse, lol

OpenStudy (anonymous):

x^2=y, integr=(1/15)log|(y-15)/y|+constant

OpenStudy (anonymous):

then you need 2x dx = dy

OpenStudy (anonymous):

oh its not so simple ,

OpenStudy (anonymous):

i.e. integral=\[\frac{1}{15}\ln \left| \frac{x^2-15}{x^2} \right|+constant\]

OpenStudy (anonymous):

when you make a substitution, you have to do dy = 2x dx

OpenStudy (anonymous):

SORRY!! out of my mind......u r right..

OpenStudy (anonymous):

then we have another problem